Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name E.Chudakov
Log entry time 00:20:08 on November 23, 2008
Entry number 250381
keyword=Experiment summary for Nov 21(swing) - 22(day)
Summary for Nov 21 (swing) - Nov 22 (day)
Smooth data taking
Beam in Use
shift |
ABU |
BANU |
BNA |
2008/11/21 swing | 6.78 |
0.08 | 1.14 |
2008/11/22 owl | 6.30 |
0.83 | 0.91 |
2008/11/22 day | 6.03 |
0.82 | 1.14 |
total | 19.01 |
1.73 | 3.19 |
Main events/issues:
- The vacuum in HRSL started deteriorating again on Friday's evening.
At 12:00 Saturday Todd took a control access and found a faulty
valve.
It is fixed and vacuum went to normal.
- The Compton group is fighting a large electronic noise in the
detector.
Plans for the next few days:
- Running with minimal changes till Tuesday;
- Tuesday: target cell change and a bunch of lesser things, to be
coordinated.